OpenStudy (anonymous):

Find the equation in standard form which passes through (–2, –7) and (1, –1). You must show all work.

OpenStudy (aroub):

-1+7/1+2 6/3 2 y=mx+b , m=2 , (1,-1) -1=2(1)+b -1=2+b -2-1=b -3=b y=2x-3